### v1

# Spectroastrometry and Reverberation Mapping of Active Galactic Nuclei. II. Measuring Geometric Distances and Black Hole Masses of Four Nearby Quasars

Abstract
The geometric distances of active galactic nuclei (AGNs) are challenging to measure because of their exceptionally compact structure yet vast cosmic distances. A combination of spectroastrometry and reverberation mapping (SARM) of broad-line regions (BLRs) constitutes a novel means to probe the geometric distance of AGNs, which has recently become practically feasible owing to successful interferometric observations with VLTI/GRAVITY. Here, we perform SARM analysis of four nearby quasars: Mrk 509, PDS 456, 3C 273, and NGC 3783. Results for the former two are reported for the first time and the latter two are revisited using our improved BLR dynamical modeling that includes the radial-dependent responsivity of BLRs. This allows us to self-consistently account for the emissivity weighting of the BLR in spectroastrometry and responsivity weighting in reverberation mapping. We obtain angular-diameter distances of the four quasars, from which we derive a Hubble constant of . Although this consititutes a large uncertainty for a measurement of , it is anticipated that the precision will improve to a competitive level once a greater number of AGNs are accessible following the upgrade of GRAVITY in the near future. From SARM analysis, the black hole masses of the four quasars are also measured with the statistical uncertainty ranging from 0.06 to 0.23 dex, consistent with the correlations between black hole masses and properties of the host bulges.
1 Introduction
Since the discovery of the first quasar, hundreds of thousands of active galactic nuclei (AGNs) have been identified from modern large-scale surveys. The popular unification model depicts an overarching generic structure of AGNs (Antonucci 1993), but many of the finer details are far from resolved. This is particularly true for the broad-line regions (BLRs) surrounding supermassive black holes (SMBHs), which are responsible for emitting broad emission lines, a definitive characteristic in the broad-band spectra of AGNs. Two observational approaches have been found to be effective and have advanced our understanding of BLR structure and kinematics. The first approach is reverberation mapping (RM; Blandford & McKee 1982; Peterson 1993) in the time domain, which can be routinely undertaken with common intermediate-aperture telescopes. The other approach probes the spatial domain with the aid of optical interferometry and is referred to as spectroastrometry (SA; Beckers 1982; Bailey 1998a; Petrov et al. 2001). It was demonstrated by the successful observation of 3C 273 (GRAVITY Collaboration et al., 2018) with the GRAVITY beam combiner mounted on the European Southern Observatory Very Large Telescope Interferometer (VLTI; GRAVITY Collaboration et al. 2017).
Over the past decades, major efforts have been made to spectroscopically monitor AGNs continually over long periods, enabling measurement of the BLR sizes through RM analysis. The typical sizes of BLRs were found to range from light-days to light-months, mainly depending on the AGN luminosity (e.g., Kaspi et al. 2000; Bentz et al. 2013; Du & Wang 2019). Such sizes are too compact to be spatially resolved by modern facilities even for the nearest AGNs. For instance, the Seyfert 1 galaxy NGC 4151 at a distance of 15.8 Mpc (Yuan et al. 2020) has a typical BLR size of light-days (Chen et al. 2023), corresponding to an angular size of milli-arcseconds. For a long time RM stood out as the most effective method to determine the spatial extent of the BLR (e.g., Kaspi et al. 2000; Peterson et al. 2004; Bentz et al. 2009; Du et al. 2014). It has also been an essential tool for the diagnostic of the gas environment in the vicinity of SMBHs (Cackett et al. 2021).
Dramatic new opportunities are now becoming available as a result of the GRAVITY instrument coming into operation. GRAVITY combines light from all four 8m Unit Telescopes of the VLTI to form six simultaneous interferometric baselines. It measures the differential phase signal of a broad emission line in the near-infrared band, which reflects SA of the emission line (namely, photocenter offset as a function of wavelength) with respect to the astrometry of the underlying continuum emission (GRAVITY Collaboration et al. 2018; Li & Wang 2023). The corresponding BLR size can then be inferred from the differential phase signal at a level of tens of micro-arcseconds. Such a high spatial resolution enables resolving the BLR in bright AGNs from the local universe (GRAVITY Collaboration et al. 2018, 2020, 2021a, 2024) up to (GRAVITY+ Collaboration et al. 2022; Abuter et al. 2024), provided these have a prominent broad emission line in the band.
Remarkably, the capabilities of SA and RM techniques are complementary and their combination (hereafter referred to as SARM) conveys a more comprehensive view of the BLR geometry and kinematics. This is reflected in the different dimensions probed by SA and RM. To be specific, SA is sensitive to the BLR structure projected onto the observer’s sky, i.e., perpendicular to the line of sight. In contrast, RM is sensitive to the structure perpendicular to the iso-delay surfaces since it measures time delays between variations of the emission line and continuum. The iso-delay surfaces have a paraboloid shape with the focus located at the central ionizing source, typically regarded as a point source coincident with the SMBH. The combination of the results from both techniques thus enable a much better overall understanding of the BLR properties as well as precise SMBH mass measurements.
Another important application of the SARM method is that it directly probes geometric distances to AGNs. The time delays measured from RM represent light travel times from the central ionizing source, typically an accretion disk, to the BLR. Multiplied by the speed of light, these thereby convey the linear size () scale of the BLR. The interferometric phases measured from SA reflect the angular astrometry of BLR photons and convey the angular size scale () of the BLR. A combination of SA and RM naturally constitutes a geometric probe for the angular-diameter distance () from the basic geometric principles (Elvis & Karovska 2002; Rakshit et al. 2015; Wang et al. 2020; Li & Wang 2023). As it is usually difficult to measure both and for the same target, extragalactic objects for which such a geometric method is applicable are exceptionally rare. The SARM approach thus offers a new, promising tool for cosmology, especially given ubiquitousness and high brightness of AGNs across the universe.
Hitherto a sample of seven nearby bright AGNs has been observed by GRAVITY (GRAVITY Collaboration et al. 2018, 2020, 2021a, 2024). Wang et al. (2020, hereafter W20) conducted the first SARM analysis on 3C 273 and GRAVITY Collaboration et al. (2021b, hereafter G21b) conducted the second application of the SARM technique on NGC 3783. These initial studies derived an angular-diameter distance Mpc with a precision of 15% for 3C 273, and Mpc with a precision of 30% for NGC 3783, respectively. They adopted a globally linear, uniform response of the BLR gas. However, there is evidence that the BLR gas might have a radial-dependent, non-linear response. The most relevant observational phenomenon comes from RM campaigns, which show that the H line widths of the mean spectra are typically broader than those of the root-mean-square (rms) spectra (see Figure 6 in Li & Wang 2025). This clearly implies that the BLR gas is more responsive at large radial distances, where the BLR gas has a relatively lower velocity (compared to the inner regions), thereby resulting in a narrower line width in the rms spectrum. Such a non-uniform response of BLRs has also been shown from the photoionization calculations (e.g., Korista & Goad 2004; Goad & Korista 2014; Zhang et al. 2021). If using the terminology “responsivity”, which quantifies the relative change of the line emissivity with respect to that of the incident continuum flux (e.g., Goad et al. 1993), the responsivity of the BLR gas is not constant, but changes with radial distance. Based on photoionization theory, the main physical reason is as follows. In the inner region where BLR gas receives a higher ionizing photon flux, the excited-state populations of hydrogen and helium atoms are enhanced, leading to a lower efficiency in converting the ionizing photons into recombination lines (such as the Balmer lines) that arise out of these excited states (Korista & Goad 2004). As a consequence, the BLR gas exhibits a low responsivity. In an extreme case where the BLR gas is fully ionized, the emission lines will no longer respond to any continuum variations, resulting in a zero responsivity. As the distance () to the ionizing source increases, the incident ionizing photon flux decreases () and the above effect gradually weakens. In other words, for an extended BLR, the changes in the responsivity of emission lines with respect to the incident ionizing photon flux naturally cause a dependence of the responsivity on radial distance.
This radially dependent responsivity of BLR gas has an important implication for SARM analysis. As SA probes emissivity weighting whereas RM probes responsivity weighting of the BLR gas, a non-uniform responsivity means that the emissivity- and responsivity-weighted BLR sizes are not the same. In this work, we perform SARM analysis on two AGNs investigated by GRAVITY Collaboration et al. (2024), Mrk 509 and PDS 456, and revisit the previous analysis on 3C 273 and NGC 3783 using the latest improved BLR dynamical modeling that includes radially dependent responsivity of the BLR gas (Li & Wang 2025). This allows us to self-consistently deal with emissivity weighting in SA and responsivity weighting in RM and therefore to alleviate inaccuracies arising from the adoption of globally uniform responses in previous analysis.
The paper is organized as follows. Section 2 compiles the observation data of SA from VLTI/GRAVITY and RM from previous optical spectral monitoring campaigns. Section 3 introduces the methodology of SARM joint analysis, including continuum light curve modeling, BLR dynamical modeling, and Bayesian inference. Section 4 summarizes estimates of the geometric distances, Hubble constant, and the black hole masses. Section 5 presents discussions on possible systematic uncertainties in the determined distances, followed by the conclusions in Section 6. Appendix A defines three types of BLR sizes, Appendix B lists the full set of model parameters and their inferred values, Appendix C shows the best fits to the differential phases of the three baselines with the strongest BLR signals, Appendix D plots the mean and rms spectra of the four AGNs, and Appendix E tests the influence of the upper limit of responsivity on the inferred model parameters.
Mrk 509 | PDS 456 | 3C 273 | NGC 3783 | |
0.0344 | 0.185 | 0.15834 | 0.009730 | |
GRAVITY Period | Jul. 2021 (2 nights) | Aug. 2018-Jul. 2021 (9 nights) | Jul. 2017-May 2018 (8 nights) | Jan. 2018-Mar. 2020 (6 nights) |
RM Period | May 2019-Nov. 2022 | May 2020-Oct. 2023 | Nov. 2008-Jul. 2018 | Feb. 2020-Jun. 2020 |
2 Observation Data
We collected the GRAVITY interferometric data and RM data of four AGNs: Mrk 509, PDS 456, 3C 273, and NGC 3783. Table 1 summarizes the observation periods (i.e., time range of the data) and the redshift used in our analysis. Below we briefly describe the main observational information as well as our additional reduction procedure to the GRAVITY interferometric data.
2.1 GRAVITY Interferometric Data
Mrk 509 was observed with GRAVITY on 25-26 July, 2021 with total on-source exposure time of 100 minutes. PDS 456 was observed during several epochs between 2018 and 2021, with total on-source exposure time of 295 minutes. The interferometric data of these two AGNs were reported in GRAVITY Collaboration et al. (2024), together with the results for two other AGNs, Mrk 1239 and IC 4329A. As these latter two are subject to large noise and the differential phase signals are less prominent, we only focus on Mrk 509 and PDS 456 in the present work. 3C 273 was observed with GRAVITY on eight nights between July 2017 and May 2018 and the interferometric data was reported in GRAVITY Collaboration et al. (2018). NGC 3783 was observed on 6 nights between January 2018 and March 2020 and the interferometric data was reported in GRAVITY Collaboration et al. (2021a).
We made additional reduction to the differential phase and line profile data of Mrk 509 and PDS 456 after applying the pipeline following the procedures described in GRAVITY Collaboration et al. (2020). We first fit and remove the static and variable instrumental effect of the differential phase for individual exposures. We also use a first-order polynomial function to remove the residual phase at the continuum emission wavelength (see Section 2.3 of GRAVITY Collaboration et al. 2020). We determine the phase error from the standard deviation of the phase at the continuum for each exposure and average all the exposures weighted according to their phase errors. To calibrate the broad Br or Pa line, we used the calibrator stars observed before or after the AGN observations to remove the instrumental and telluric features. Thereafter, we apply a third-order polynomial function to fit the calibrated continuum and normalize the line profile to the continuum. For Mrk 509 and PDS 456, we prefer using early-type stars to calibrate the spectra because they have less absorption features, with only the Br absorption line significant in the relevant spectral range, which imprints a narrow Br emission line on the final calibrated AGN spectra. To subtract this contaminating component, we use two Gaussian functions to fit the AGN Br or Pa line and a Lorentzian function to fit the narrow Br line. The AGN and stellar Br lines are not strongly blended (if at all) because the AGN Br line is significantly redshifted. We notice that, compared to the double-Gaussian function used to fit the stellar Br line as by GRAVITY Collaboration et al. (2024), the Lorentzian function can improve the fit to the Pa line profile of PDS 456. This is because the double-Gaussian function for the stellar Br line cannot be properly separated from those for the AGN Pa line, thereby causing an inappropriate subtraction around the blue wing of the AGN line. We directly adopt the data of 3C 273 and NGC 3783 from GRAVITY Collaboration et al. (2018) and GRAVITY Collaboration et al. (2021a), respectively. The NGC 3783 data were reduced in the same way as described above. The 3C 273 data were however reduced differently. Instead of decomposing the instrumental differential phase signatures into static and variable components, GRAVITY Collaboration et al. (2018) incorporated a Gaussian kernel with an FWHM of 24 pixels to convolve the differential phase spectra and generate the broad instrumental differential phase. We confirmed that it makes little difference if we conduct the joint analysis with the data reduced with the new approach.
2.2 Reverberation Mapping Data
Li et al. (2024) reported multi-year RM data of five AGNs (IC 4329A, Mrk 335, Mrk 509, Mrk 1239, and PDS 456), selected from the sample of potential AGNs suited for GRAVITY/GRAVITY+ observations compiled by W20. For Mrk 509 and PDS 456, the optical spectroscopic and photometric monitoring were conducted from May 2019 to November 2022, and from May 2020 to October 2023, respectively, with a typical sampling cadence of every 4-5 days. The photometric data were further combined with the public archival data from the Zwicky Transient Facility (ZTF; Graham et al. 2019) and the All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al. 2014; Kochanek et al. 2017) to yield high-cadence continuum light curves. These different sources of data were intercalibrated with the Bayesian software PyCALI111https://github.com/LiyrAstroph/%****␣ms_v5.tex␣Line␣225␣****PyCALI (Li et al. 2014), which uses a multiplicative factor and an additive factor to account for the inhomogeneous apertures and differences in flux calibration. For Mrk 509, from the four years of monitoring, the H time delay (around 50 days) can be detected in three years (2019, 2021, and 2022). However, when using the full data span, there appears a long-term trend in the H variations, leading to an apparently long time delay component (around 300 days). The explanation for such a long-term trend is still under debate. We therefore only use one-year data for SARM analysis. We choose the RM data for year 2022 since it shows the most significant variability. We also present analysis with the RM data in 2019 and 2021 and find that the obtained distances are consistent with each other within uncertainties. For PDS 456, the H time delay is about 250 days and we use the four-year RM data for the SARM analysis.
For 3C 273, Zhang et al. (2019) presented the analysis of 10 years of optical monitoring data between November 2008 and July 2018, using the public archival data from the Steward Observatory spectropolarimetric monitoring project (Smith et al. 2009) and new observations from the super-Eddington accreting massive black hole project (e.g., Du et al. 2014). 3C 273 is radio loud and the optical emission from the beamed jet is not negligible compared to that from the accretion disk. As demonstrated by Li et al. (2020), the jet contamination results in a distinct long-term component in the continuum light curve that does not have a corresponding echo in the H fluxes. Such jet contamination can be partly corrected by subtracting a linear long-term trend. As we show below, we set the slope of the linear trend as a free parameter to be determined from the Bayesian inference.
The latest published RM study of NGC 3783 was conducted by Bentz et al. (2021) throughout the first half of 2020 using the Las Cumbres Observatory global telescope network. A total of 50 spectra were obtained with a typical sampling interval of 2 days. The 5100 Å data was merged with the -band photometric data (with 209 epochs), resulting in a continuum light curve spanning a time span of 137 days from JD 2,458,891 to 2,459,028 (see Figure 4 in Bentz et al. 2021). NGC 3783 has also been covered by the ASAS-SN survey since February 2012. We use the PyCALI software to scale and shift the ASAS-SN light curve to match the merged light curve of Bentz et al. (2021). This extends the temporal baseline of the continuum light curve from 137 days to 268 days in the 2020 season and also supplies continuum variation information for other seasons. The former is useful to constrain the BLR model and the latter helps to better characterize the continuum variability.
In modeling RM data, we need to subtract the components in the continuum light curves irrelevant to emission-line reverberation. One such component is the host galaxy’s starlight. Among the four AGNs, Mrk 509, PDS 456, and 3C 273 have a high optical luminosity at 5100 Å (Zhang et al. 2019; Li et al. 2024) and the host galaxy contribution can be treated as negligible. This is supported by the spectral decomposition, which did not identify a significant host galaxy component (Zhang et al. 2019; Li et al. 2024). For NGC 3783, we subtract the host galaxy component () according to the HST image decomposition employed by Bentz et al. (2021). 3C 273 is additionally subject to the contamination from the jet emission. We subtract an averaged jet flux from the continuum light curve according to the decomposition of the jet and accretion disk emissions proposed by Li et al. (2020), which derived an averaged jet flux to the total continuum flux fraction of .
As summarized in Table 1, we note that the GRAVITY and RM observations of all the four objects are undertaken in overlapping periods. This largely minimizes the influences of possible secular changes in BLR kinematics, which occur in the dynamical timescale, typically of the order of several years or longer.
3 SARM Joint Analysis
The procedure underpinning the SARM joint analysis has been described in a number of previous studies, such as W20, G21b, Li et al. (2022), and Li et al. (2024). What distinguishes our analysis from these earlier studies is that we include the latest improvement of BLR dynamical modeling that takes into account radial-dependent responsivity of the BLR (see Li & Wang 2025 for details). This improvement allows us to accommodate for responsivity weighting in SA and emissivity weighting in RM (e.g., Zhang et al. 2021; Li et al. 2022; Li & Wang 2025). For the sake of completeness, we briefly describe the major points of SARM analysis below.
To correct for the cosmological time dilation, we reduce the observation times of the RM light curves by a factor of , where is the redshift of the object. Since the RM analysis depends only on the time differences between data points, this manipulation equivalently converts the light curves into the rest frame.
3.1 Continuum Light Curve Modeling
We first need to evaluate the continuum fluxes on a sufficiently well sampled time grid, which is required to integrate reprocessed emission from the whole BLR with a wide range of time delays. To this end, we employ the damped random walk (DRW) model to delineate the continuum variability, which is a stationary Gaussian process with well-established procedures to interpolate and extrapolate the observed continuum light curve (e.g., Rybicki & Press 1992; Foreman-Mackey et al. 2017; Li et al. 2018). We use a series of evenly spaced points to represent the continuum reconstruction. This series needs to cover an adequate time period prior to the first epoch of the emission line so as to appropriately model the emission-line flux at any given observation epoch. Specifically, the time length of this period should be equal to or longer than the maximum time delay of the BLR. These continuum points plus two hyper-parameters of the DRW process (the damping time scale and variation amplitude) are treated as free parameters and their best-fit values are determined during the subsequent Bayesian inference (see below).
3.2 BLR Dynamical Modeling
We construct a BLR dynamical model following Li & Wang (2025), which made improvements to the previous modeling scheme of Pancoast et al. (2014) and Li et al. (2018) by including radial-dependent responses. The BLR is assumed to be composed of a large number of point-like, non-interacting clouds, moving in the gravitational well of the central SMBH. These clouds reprocess the ionizing continuum emissions into line emissions, with variable emissivity and responsivity depending on the clouds’ radial location. Here, responsivity quantifies the relative change of line emissivity with respect to that of the incident continuum222Responsivity is mathematically defined as , where is the cloud’s emissivity and is the ionizing flux illuminating the cloud. (Korista & Goad 2004). The optical continuum is used as a surrogate to the unobservable ionizing continuum as their inter-band time delay is negligible compared to the typical time delays of the BLR.
BLR clouds are distributed within a thick disk, with an opening angle and inclination angle (see Figure 1 of Li et al. 2013). The radial distribution of the clouds follows a shifted gamma distribution, namely, the cloud’s radial distance from the central black hole is given by , where is the Schwarzschild radius, is the mean BLR radius, is the fraction of the inner BLR edge relative to , and is a random number drawn from a gamma distribution with and . The parameter quantifies the standard deviation of the BLR cloud’s radial distance from the central point in units of . It also controls the shape of the radial distribution, which may be made to be exponential (), long-tailed (), or narrow peaked (). Within the opening angle, BLR clouds are distributed uniformly in the azimuthal -direction. In the vertical -direction, the angular dispacement of a cloud is assigned as , where the parameter controls the extent to which clouds are clustered over the outer face of the BLR disk and is a random number drawn from a uniform distribution between 0 and 1 (Pancoast et al. 2014).
The motion of BLR clouds is parameterized to account for various kinematics, including elliptical Keplerian orbiting and bound and unbound inflow/outflow (Pancoast et al. 2014). A fraction of clouds are in elliptical Keplerian orbits and the remaining fraction of clouds are either inflowing or outflowing. The latter is controlled by a parameter , with indicating inflow and indicating outflow. The velocity of each cloud is first determined for the orbital plane in which it moves around the central SMBH and then converted into real three-dimensional velocity through coordinate transformation. For elliptical orbits, radial and tangential velocities in the plane are drawn from Gaussian distributions centered around the point along the ellipse connecting circular orbiting velocity () and radial escape velocity (). For inflowing/outflowing clouds, velocities are assigned similarly, except that the Gaussians are centered around . In addition, the Gaussian centers are allowed to shift along the ellipse, controlled by a parameter , to mimic the possibility that clouds are inflowing or outflowing at a velocity smaller than the escape velocity (see Pancoast et al. 2014). It is important to note that the SA signal is sensitive to the degree of coherent rotation of the BLR clouds. All clouds’ tangential velocities () are forced to have the same sign (e.g., positive) to ensure they rotate coherently.
From photoionization calculations (Korista & Goad 2004), it is known that the reprocessing coefficient of BLR clouds, denoted by , has a radial dependence. The reprocessing coefficient together with the radial distribution of clouds determine the broad-line emissions. To reduce the degrees of freedom, we let be absorbed into the parameterization of the radial distribution and simply set for all clouds. This means that the radial distribution of BLR clouds includes emissivity weights and should not simply be treated as the number density distribution.
We now introduce a parameterization of the responsivity of clouds as (Li & Wang 2025)
where , , and are free parameters. This form of responsivity can approximately reproduce the photoionization calculations in which the BLR clouds are more responsive at large radius (e.g., Korista & Goad 2004; Li & Wang 2025). In our calculations, the responsivity is further limited to a range and those outside this range are forced to take the lower/upper limits. The lower limit is set to ensure positive responses of BLR clouds to the continuum variations, while the upper limit is set to align with previous photoionization calculations, which generally found (e.g., Korista & Goad 2004; Goad & Korista 2014, 2015; Zhang et al. 2021; Li & Wang 2025). We also include the following two anisotropic effects. Each cloud is assigned an emission weight , where is a free parameter and is the angle between the observer’s and cloud’s line of sight to the central continuum source. In addition, the transparency of the BLR equatorial plane is controlled by a parameter , which represents the fraction of clouds below the equatorial plane that is not obscured and contributes to the observed line emissions.
With the application of the above BLR model, the emissivity- and responsivity-weighted transfer functions at time delay and velocity are obtained by summing over all clouds (Li & Wang 2025)
where is the Dirac delta function and the subscript refers to the -th cloud.
3.3 Calculating Flux, Profile, and Spectroastrometry of the Emission Line
By recasting the continuum into constant and variable components , the emission line flux is calculated as (Li & Wang 2025)
where and are velocity integrals of the transfer functions calculated by Equations (2) and (3). On the right-hand side of Equation (4), the first term represents the constant component and the second represents the variable component of the emission line light curve. The emission-line profile is given by
where is the delay integral of the emissivity-weighted transfer function.
The differential phase at velocity is calculated as
where is the flux ratio of the emission line relative to the underlying continuum, is coordinate of the baseline, is the emissivity-weighted offset of the BLR projected onto the sky, is the offset of the continuum emission, and is the angular-diameter distance. The offset is determined by summing over all clouds as
where is the -th cloud’s offset, is the -th cloud’s three-dimensional coordinate, and is the unit vector in the line of sight.
It is worth reiterating that because of the radial-dependent responsivity, the emissivity- and responsivity-weighted BLR sizes are no longer the same (see Appendix A for the definitions). For a radially increasing responsivity ( and ), the BLR clouds are more responsive at larger radius. This causes the responsivity-weighted size obtained from RM to be larger than the emissivity-weighted size obtained from SA (Li & Wang 2025). The inclusion of responsivity in dynamical modeling enables us to self-consistently combine these two observational techniques and eliminate any potential biases that may arise from imposing a uniform BLR response in SARM analysis.
3.4 Bayesian Inference
Given the above continuum light curve and BLR dynamical modeling, we can calculate the modeled SA and RM data and compare these against observations to determine the posterior distributions of model parameters using the Markov Chain Monte Carlo (MCMC) technique. The SA data include the profile and differential phases of the broad Pa or Br line. The corresponding likelihood function is written
where the subscript represents -th velocity bin, the superscript represents the modeled quantity, and are the differential phase and uncertainty, and and are the line flux density and uncertainty, respectively. The RM data include the H and continuum light curves. The corresponding likelihood function is given by
where the subscript represents -th epoch, the superscript represents the modeled quantity, and and are the H line flux and uncertainty, respectively. Here, the modeled line flux depends on the continuum light curve modeling.
The joint likelihood function of SARM analysis is the product of Equations (8) and (9), namely, . The posterior probability distribution of model parameters is then , where is the parameter prior. In Appendix B, we list the full model parameters and their priors. We employ the diffusive nested sampling (DNS) algorithm (Brewer et al. 2011) to explore the parameter space and generate Markov chains, which has been shown to be effective at exploring multimodal distributions and strong correlations between parameters. We have incorporated the above BLR dynamical modeling into our previously developed package BRAINS333The latest code is publicly available at https://github.com/LiyrAstroph/BRAINS, while the version used in this work is available at https://doi.org/10.5281/zenodo.10700132. (Li et al. 2013, 2018), which utilizes the DNS code CDNest444The latest code is publicly available at https://github.com/LiyrAstroph/CDNest, while the version used in this work is available at https://doi.org/10.5281/zenodo.10700132.. Both BRAINS and CDNest are written in C Language with the standardized message passing interface and are thereby portable to a wide range of supercomputer clusters. From the generated posterior samples of parameters, we assign the best estimates by the median values and the uncertainties by the 68.3% confidence interval.
4 Results
4.1 Overview
In Figures 1-4, we show the SARM analysis results of the four AGNs, including fits to the SA and RM data, and posterior distributions of the BLR size, black hole mass, and angular-diameter distance. Table 2 lists the best estimates and uncertainties of these three key parameters, together with the BLR angular size. A full list of the inferred parameter values and uncertainties are summarized in Appendix B. For the sake of clarity, we only plot the differential phases averaged over the three baselines with the strongest BLR phase signals. The continuum phases have also been subtracted. The best fits to the individual differential phases of these three baselines are presented in Appendix C.
When running the MCMC, we find that for PDS 456, the continuum phase offset parameter is strongly degenerate with the angular-diameter distance. The Bayesian modeling always tends to fit the whole differential phases with the continuum phase, leaving quite small phase residuals for the BLR emissions. This in turn results in an unrealistically large angular-diameter distance. We therefore switch off the continuum phase offset parameters for PDS 456. As illustrated below, the distance obtained without including the continuum phase offset turns out to be more reasonable. In addition, we also note that the H light curve, shown in the left bottom panel of Figure 2, displays some fine variation patterns that do not have counterparts in the continuum light curve. It seems that these non-echoing H variations are not solely due to noise but their origin is unclear. The analysis reproduces only the averaged variation trend of the H light curve. As a result, there appear quite a few data points deviating from the best reconstruction.
In the averaged differential phase curves (with the continuum phase subtracted) plotted in the second panel of the top left row of Figures 1-4, 3C 273 and NGC 3783 display an approximately S-shaped signal typical for Keplerian rotation, which is characterized by a crest and trough respectively on either side of the center wavelength of the emission line. In contrast, PDS 456 and Mrk 509 show an asymmetric signal dominated by a trough in the differential phase curve, indicative of non-Keplerian motion. In the present dynamical modeling, the latter can be explained by the presence of a fraction of BLR clouds with inflowing/outflowing motion plus anisotropic emissions.
4.2 The Inferred Radial-dependent Responsivity
Figure 5 illustrates the obtained responsivity as a function of radius, along with the 1 confidence intervals. We find that 3C 273 and PDS 456 show a strong radial-dependent responsivity. In contrast, the dependence for Mrk 509 and NGC 3783 is relatively weak. For a strongly radially increasing responsivity, we expect a significant difference in the emission-line widths from the mean and rms spectra in the RM data (see Li & Wang 2025). This is because the BLR clouds are more responsive at greater radii with a lower velocity. These clouds therefore add larger weights to variations of the line flux density at low-velocity bins, resulting in a narrower line width in rms spectra compared to the mean spectra. In earlier studies in literature, two types of line width are used, the full width at half maximum (FWHM) and the line dispersion (). In Appendix D, we show the mean and rms spectra from previous RM campaigns for the four AGNs. For 3C 273, the H FWHM in the rms spectrum () is narrower than in the mean spectrum () by a factor of 0.4 (Zhang et al. 2019). The same is the case for the H . Similar to 3C 273, PDS 456 also has significantly narrower H line widths in the rms spectrum (Li et al. 2024). In contrast, the H line widths of Mrk 509 (Li et al. 2024) and NGC 3783 (Bentz et al. 2021) are similar for the mean and rms spectra. Our obtained responsivities in Figure 5 are remarkably coincident with the difference between the mean and rms H line widths, implying a validation of our dynamical modeling. The responsivity can also be further tested by future photoionization calculations.
The responsivity directly affects the variation amplitude of the emission line. From Equation (4), we can deduce that the responsivity tends to scale up () or down () the variation amplitude of the emission line relative to that of the continuum. As mentioned in Section 2.2, in real observational data the continuum flux also contains the extra contribution from the host galaxy’s starlight or other components, causing contamination to the constant continuum component. This in turn requires a larger responsivity to reproduce the fractional variability amplitude of the emission line. As a result, the inferred responsivity will be biased to a larger value in the presence of other contamination. In our analysis, we have corrected for the contamination by the host galaxy for NGC 3783 and the jet emission for 3C 273. The contamination of the host galaxy for the other two objects Mrk 509 and PDS 456 may be treated as negligible. We also run SARM analysis with the original uncorrected RM data and confirm that these corrections do not affect the obtained geometric distances.
In our analysis, we set an upper limit of the responsivity of (see Equation 1), based on previous photoionization calculations (e.g., Korista & Goad 2004; Goad & Korista 2014, 2015; Zhang et al. 2021; Li & Wang 2025). In Appendix E, we rerun analysis with two different upper limits of and and illustrate the influences on the BLR size, SMBH mass, and angular-diameter distance. We find that the results for Mrk 509, PDS 456, and NGC 3783 are quite insensitive to the choice of . For 3C 273, the best estimates of the BLR size and angular-diameter distance are relatively smaller for , but still consistent with those of and within 1 uncertainties.
Parameter | Mrk 509 | PDS 456 | 3C 273 | NGC 3783 |
Note. — The BLR angular size is defined by .
4.3 Geometric Distances
The geometric distances of Mrk 509 and PDS 456 are measured here for the first time, while the analysis of the other two sources, 3C 273 and NGC 3783, had been conducted by W20 and G21b, respectively, based on linear BLR responses. Below we first present our distance measurements of Mrk 509 and PDS 456, and then our latest measurements of 3C 273 and NGC 3783, together with a detailed comparison with the previous results. The left panel of Figure 6 shows the distribution of the four quasars in the plane of redshift versus angular-diameter distance.
We obtain a distance of for Mrk 509 and for PDS 456. For comparison, if using a flat CDM cosmology with , , and , the corresponding angular-diameter distances are and 2.83, respectively. Those values are compatible with our inferences within 1 uncertainties.
Our inferred angular-diameter distance of 3C 273 is and of NGC 3783 is . The corresponding distances estimated from the cosmology are 2.77 and 1.63, respectively, again compatible with our inferences within 2 uncertainties. NGC 3783 is quite nearby and therefore subject to non-negligible peculiar motion (see G21b for a detailed investigation). If using a peculiar velocity (see the next section for details), the corrected redshift will be smaller by about 5%. The anticipated for NGC 3783 from cosmology will be reduced by the same factor, to . In any case, our obtained geometric distances are generally consistent with the cosmology-based estimates.
Using SARM analysis, W20 derived for 3C 273 and G21b derived for NGC 3783. Considering the associated uncertainties of the inferred distances, these values are consistent with our results. However, we stress that the BLR dynamical modeling in both W20 and G21b adopted a globally linear, uniform responsivity, which always results in the same emissivity- and responsivity-weighted BLR sizes (see also Li & Wang 2025). Our improved SARM analysis naturally accounts for the differences in these two BLR sizes and largely alleviates the increased systematic biases in distance measurements. As mentioned in the preceding section, we find that 3C 273 shows strong radial dependence of the responsivity, while this dependence is weak in NGC 3783 (see Figure 5). These results remarkably coincide with the line widths in the mean and rms spectra of the two objects (see Appendix D).
For 3C 273, the smaller center value of by 0.16 dex we obtained compared to the one given by W20 can be understood as follows. Due to the use of linear responses in W20, the emissivity-weighted BLR size of 3C 273 tends to be overestimated and the resulting distance is therefore forced to be larger given the same differential phase data. For NGC 3783, G21b demonstrated that an inclusion of non-linear responses (still globally uniform) results in a distance about 10% larger. Note that we included the photometric light curve from the ASAS-SN archival data, which significantly extends the temporal baseline of the continuum light curve and therefore enables better continuum modeling. For example, the spectral monitoring of Bentz et al. (2021) started on JD2,458,891, while ASAS-SN database includes -band photometry since JD2,458,812, which is 79 days earlier (see Figure 4). The above two factors might explain the reason why our inferred distance is larger by 0.1 dex than that of G21b, although both are still consistent within uncertainties.
The statistical uncertainties in of the four objects range from 0.09 to 0.24 dex, stemming from the uncertainty of the linear size (constrained by the RM data) and angular size (constrained by the SA data) of the BLR. The inferred values in Table 2 imply that the uncertainties of mainly come from the SA data for 3C 273 and NGC 3783, whereas for Mrk 509 and PDS 456, the RM and SA data make comparable contributions to the uncertainties.
4.4 An Estimate of the Hubble Constant
The measured angular-diameter distances allow us to estimate the Hubble constant. For a flat CDM cosmology, the relation between redshift and angular-diameter distance is
where is the speed of light, is the Hubble constant, is the dimensionless density of mass, and the dimensionless density of dark energy . Because of the rather limited redshift span of the four sources, the parameter can not be constrained properly and we simply adopt . Note that this value has little effect on the determination of the Hubble constant. For NGC 3783, according to the comprehensive investigation by G21b, the peculiar velocity was estimated to be from the averaged peculiar velocity of 11 galaxies within 8 Mpc centered on the position of NGC 3783 obtained from the 6dF galaxy redshift survey peculiar velocity map (Springob et al. 2014). We correct the redshift of NGC 3783 by taking into account the peculiar velocity as (Davis & Scrimgeour 2014)
where is the normal recession redshift. This leads to .
By fitting the inferred angular-diameter distances of four AGNs with Equation (10), we obtain a Hubble constant of . In the right panel of Figure 6, we plot the posterior distribution of , along with the measurements from the Planck satellite (Planck Collaboration et al. 2020) and SH0ES project (Reid et al. 2019). Our estimated Hubble constant bears a large uncertainty (15%) that is insufficiently precise to address the current Hubble tension problem (e.g., Riess 2020). However, the forthcoming upgraded instrument GRAVITY+ will significantly improve the sensitivity magnitude down to , allowing observations of a few hundred AGNs with higher-quality differential phase measurements (GRAVITY+ Collaboration et al. 2022), which will enable the SARM approach to achieve a competitive percent-level precision (Songsheng et al. 2021).
4.5 BLR Sizes and SMBH Masses
As mentioned above, the radial-dependent responsivity results in different responsivity- and emissivity-weighted BLR sizes (denoted as and , respectively). Moreover, owing to the inclusion of anisotropic effects of BLR emissions (see Section 3.2), the centroid time delay of the responsivity-weighted transfer function , which is a quantity directly related to RM analysis, might not be exactly the same as . In Appendix A, we illustrate how to calculate these BLR sizes given a BLR dynamical model. Figure 7 shows comparisons among the obtained BLR sizes for the four sources. As expected, is larger than by for PDS 456 and 3C 273, both showing a significant radially increasing responsivity (see Figure 5). There appears a good consistency between and because the transfer function is closely related to the responsivity as illustrated in Equation (3). As a consequence, the size of both PDS 456 and 3C 273 is smaller than by a factor of 60%. These results imply that when comparing the BLR sizes obtained from interferometry (corresponding to ) against those from RM (corresponding to ), one needs to carefully take into account the differences in the BLR dimensions represented by and . In other words, Figure 7 by no means indicates that GRAVITY interferometry underestimates the BLR sizes or RM overestimates the BLR sizes. Indeed, these two approaches probe different BLR measures. To further illustrate the above point, we apply the same BLR dynamical model to only GRAVITY data and make a comparison of the obtained angular sizes () with the sizes from our SARM analysis in Figure 8. Here, is used to calculate the angular sizes in SARM analysis. We can find a good agreement between the two approaches. This in turn means that when using to calculate the SARM angular size, there will be a significant difference from the SA angular size for PDS 456 and 3C 273.
For 3C 273, GRAVITY Collaboration et al. (2018) measured a BLR angular size of as, consistent with our result of as (see Table 2). However, we note that the linear BLR size of ld reported by GRAVITY Collaboration et al. (2018) is larger than our inference of ld. This is because they adopted a distance of 550 Mpc, larger than our measured distance of Mpc. For the other three sources, we can also find an agreement with previous SA analysis within uncertainties (GRAVITY Collaboration et al. 2021a, 2024).
We obtain a black hole mass for Mrk 509, for PDS 456, for 3C 273, and for NGC 3783. Recently, using high-spectral-resolution integral-field spectroscopic observations, Winkel et al. (2024) measured the stellar velocity dispersions and dynamical masses of the host bulges for a sample of nearby AGNs, including Mrk 509, 3C 273, and NGC 3783. In Figure 9, combined with our obtained SMBH masses, we superimpose these three objects on the scaling correlations between SMBHs and bulges compiled by Kormendy & Ho (2013). We find that the three objects match these correlations within the associated uncertainties.
By comparing with SMBH mass determinations using only RM, our inferred masses, except for PDS 456, are generally consistent with previous RM results within uncertainties (Zhang et al. 2019; Bentz et al. 2021; Li et al. 2024). When using a unity virial factor and the H FWHM from the mean spectrum, the RM campaign of Li et al. (2024) reported a mass of for PDS 456. Note that such a RM-based mass estimate does not include the scatter of the virial factor (0.3 dex; e.g., Ho & Kim 2014), hence underestimating the uncertainties. The optical spectrum of PDS 456 shows strong iron emissions and the [O III]4959, 5007 lines are almost invisible (see Appendix D). These features are commonly seen in narrow-line Seyfert I galaxies and are believed to be indicative of super-Eddington accretion (e.g., Boroson & Green 1992; Collin & Kawaguchi 2004; Du et al. 2014). Thus a lower SMBH mass might be more plausible. Future comprehensive observations with improved data quality in both SA and RM are highly worthwhile to verify the present black hole mass estimates.
Compared to the SMBH mass inferred from the GRAVITY data alone (GRAVITY Collaboration et al. 2018, 2021a, 2024), we again find overall consistency within 1 uncertainties. In particular for PDS 456, the SA analysis by GRAVITY Collaboration et al. (2024) obtained an SMBH mass of . Despite this value differing from our result by 0.47 dex, the lower error of dex implies that the two results are statistically in agreement.
5 Discussion
5.1 Differences between the H and Pa/Br Lines
An implicit assumption in the present SARM analysis is that the optical H line probed by RM and the near-infrared Pa/Br line probed by SA stem from exactly the same region. In practice, the observed profiles of these lines are slightly different. Figure 10 compares the Pa/Br line with the H line subtracted from the mean spectra of the RM campaign, all normalized by the peak flux density. The Pa/Br lines are also manually convolved with a Gaussian to match the spectral resolution of the H lines. As observed, the Pa/Br line is overall narrower than H. After subtracting line width broadening arising from the spectral resolution, we estimate the line widths as follows: FWHM(H, Pa)=(2840, 2620) km s-1 for PDS 456, FWHM(H, Pa)=(3220, 2580) km s-1 for 3C 273, FWHM(H, Br)=(3650, 3560) km s-1 for Mrk 509, and FWHM(H, Br)=(4340, 3680) km s-1 for NGC 3783 (see also GRAVITY Collaboration et al. 2021a). The resulting line width ratio between the H and Pa/Br lines is 1.09, 1.25, 1.02, and 1.18, respectively. Using the virial relation , the corresponding differences in the BLR sizes of the two emission lines become 18%, 56%, 5%, and 39%, respectively.
The fact that the H line is relatively broader means that current SARM analysis underestimates the geometric distance. If simply scaling results with the above estimated differences in the BLR sizes, the geometric distance will increase by roughly 0.07, 0.19, 0.02, and 0.14 dex for PDS 456, 3C 273, Mrk 509, and NGC 3783, respectively. However, we note that a precise evaluation of the influences to the distance measurements requires comprehensive modeling of the H and Pa/Br lines simultaneously, which is beyond the scope of the present work. Nevertheless, our estimate of the Hubble constant remains within a reasonable range, indicating that the discrepancy arising from different emission lines measured by SA and RM is moderate. In future, this issue can be surmounted by near-infrared RM campaigns observing the same broad emission line as SA or a joint analysis of SA and velocity-resolved RM. In the latter approach, one no longer needs to assume the same region for the H and Pa/Br lines but only let the two lines share the same central SMBH (Li et al. 2022).
5.2 Results of Mrk 509 from Different Years of RM Data
Mrk 509 was monitored for four years between 2019 and 2022 and the H time delay can be detected in three years, except for 2020 due to the absence of clear variation patterns in the RM data (Li et al. 2024). We choose the 2022 RM data for SARM analysis given that this year marked the largest variability. In Figure 11, we illustrate a comparison of the obtained BLR size, SMBH mass, and angular-diameter distance between years 2019, 2021, and 2022. As observed, the best estimates are consistent with each other within uncertainties, although the BLR size in 2021 shows a slightly larger median value. This is as expected since the H time delays in these three years are in overall agreement.
6 Conclusions
We combine SA and RM data to perform joint SARM analysis to measure the geometric distances of four AGNs: Mrk 509, PDS 456, 3C 273, and NGC 3783. SARM results are reported for the first time for the former two sources and are revisited for the latter two. Inspired by photoionization calculations and the long-standing observation from RM campaigns that H emission-line widths in rms spectra are generally narrower than those in mean spectra, we improve BLR dynamical modeling by including a radially dependent responsivity. This enables a self-consistent treatment of the emissivity weighting of the BLR in SA and responsivity weighting in RM and thereby eliminates any potential biases that may arise from imposing a uniform, linear BLR response. Such a non-linear, radially dependent responsivity has been neglected in all previous dynamical modeling approaches. We obtain an angular-diameter distance of for Mrk 509, for PDS 456, for 3C 273, and for NGC 3783, from which we derive a Hubble constant of . The SARM analysis also allow us to measure a black hole mass for Mrk 509, for PDS 456, for 3C 273, and for NGC 3783. These values are in good agreement with known relations between black hole masses and stellar velocity dispersions or dynamical masses of the host bulges (though for PDS 456 without relevant measurements of the host galaxy).
Although our estimated Hubble constant still carries a large uncertainty, the forthcoming upgraded instrument GRAVITY+ will be able to observe more AGNs (up to a few hundred) with higher-quality differential phase measurements (GRAVITY+ Collaboration et al. 2022). A sizeable RM campaign coordinated with GRAVITY+ interferometric observations will enable the SARM approach to achieve a competitive precision for the Hubble constant. In light of the geometric nature of the measured distance, the SARM approach will ultimately offer a promising cosmic probe, independent of all other cosmic tools.
Acknowledgements
We thank Misty Bentz for kindly sharing the RM light-curve data of NGC 3783 and Hannah Übler for useful comments on the manuscript. YRL acknowledges financial support from the National Key R&D Program of China (2021YFA1600404, 2023YFA1607904), the National Natural Science Foundation of China (12273041), and from the Youth Innovation Promotion Association CAS. JMW acknowledges financial support from the National Natural Science Foundation of China (11991050, 12333003). LCH was supported by the National Key R&D Program of China (2022YFF0503401), the National Science Foundation of China (11991052, 12233001), and the China Manned Space Project (CMS-CSST-2021-A04, CMS-CSST-2021-A06).
We thank the support of the staff of the Lijiang 2.4 m telescope and of the CAHA 2.2m telescope. This work is based on observations collected at the Centro Astronómico Hispanoen Andalucía (CAHA) at Calar Alto, operated jointly by the Andalusian Universities and the Instituto de Astrofísica de Andalucía (CSIC). This work has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.
Appendix A Calculating the BLR Sizes
From the transfer function in Equation (3), the BLR size from the centroid time delay is calculated as
which can be recast into
The responsivity-weighted size is calculated by summing up all clouds’ radii with weights as (Li & Wang 2025)
Similarly, the emissivity-weighted size is obtained by
which is indeed equal to the parameter because the weights do not depend on radius. It is easy to show that will be exactly equal to only if is a global constant.
Appendix B The Full List of Model Parameters and Their Best Estimates
Appendix C Full Fits to Spectroastrometric Data
In Figure 12, we show the best fits to the differential phases of the three baselines with the strongest BLR signals for the four AGNs. The differential phases are averaged over epochs.
Appendix D The Optical Spectra
For the sake of completeness, we show the optical mean and rms spectra of Mrk 509, PDS 456, 3C 273, and NGC 3783 from previous RM observations in Figure 13. As observed, PDS 456 and 3C 273 show weak [O III] and strong Fe II emissions, which are typical features of super-Eddington AGNs (e.g., Boroson & Green 1992; Collin & Kawaguchi 2004; Du et al. 2014).
Appendix E The Upper Limit of the Responsivity
In our calculations, the responsivity is parameterized as a power law of radial distance (see Equation 1). The maximum responsivity is limited to according to previous photoionization calculations (e.g., Korista & Goad 2004; Goad & Korista 2014, 2015; Zhang et al. 2021; Li et al. 2024). To test whether this limit affects the parameter inference, we rerun the SARM analysis using an upper limit of and . In Figure 14, we show a comparison of the obtained BLR size, SMBH mass, and angular-diameter distance for these different upper limits. We find that for Mrk 509, PDS 456, and NGC 3783, the results are quite insensitive to the choice of . For 3C 273, the upper limit of tends to yield a smaller BLR size and thereby a smaller angular-diameter distance. Nevertheless, the differences from those of and are about 0.1 dex, still within 1 uncertainties, implying that the influence of on the distance measurement is not significant. A BLR dynamical model with an inclusion of photoionization models will relax the need of present simple parameterization for the responsivity and help further improve SARM analysis. This is beyond the scope of this paper and will be deferred to a future work.
Parameter | Prior | Range | Note |
(ld) | LogUniform | (1, /2) | Emissivity-weighted mean BLR size |
Uniform | (0, 1) | Fraction of the inner BLR edge in unit of | |
Uniform | (0, 2) | Shape parameter of radial distribution of BLR clouds | |
(∘) | CosUniform | (0, 90) | Inclination angle |
(∘) | Uniform | (0, 90) | Openning angle |
Uniform | (-1/2, 1/2) | Azimuthal anisotropic parameter of cloud emission | |
Uniform | (0, 5) | Clustring parameter of BLR clouds towards the surfaces | |
Uniform | (0, 1) | Transparency of the BLR equatorial plane | |
Uniform | (0, 1.5) | Constant coefficient of responsivity | |
Uniform | (0, 1.5) | Power-law amplitude of responsivity | |
Uniform | (0, 5) | Power-law index of responsivity | |
PA (∘ E of N) | Uniform | (0, 360) | Position angle |
(ld) | Uniform | (0, 100) | Longitudinal offset of the continuum emission projected on the sky |
(ld) | Uniform | (0, 100) | Latitudinal offset of the continuum emission projected on the sky |
(Mpc) | LogUniform | (10, 104) | Angular-diameter distance |
() | LogUniform | (106, 109) | Black hole mass |
Uniform | (0, 1) | Fraction of clouds in elliptical orbits | |
Uniform | (0, 1) | Flag for determining inflowing or outflow orbits | |
(∘) | Uniform | (0, 90) | Rotation angle of inflow or outflow orbits |
Uniform | (0, 2) | Constant component of the continuum light curve |
Note. — denotes the time span of the emission line light curve. “LogUniform” denotes a uniform prior distribution for the logarithm of the parameter and “CosUniform” denote a uniform prior distribution for the cosine of the parameter.
Parameter | Mrk 509 | PDS 456 | 3C 273 | NGC 3783 |
PA (∘ E of N) | ||||
(ld) | 0 | |||
(ld) | 0 | |||

### v2

@@
 We obtain angular-diameter distances of the four quasars, from which we derive a Hubble constant of .
-Although this consititutes a large uncertainty for a measurement of , it is anticipated that the precision will improve to a competitive level once a greater number of AGNs are accessible following the upgrade of GRAVITY in the near future.
+Although this constitutes a large uncertainty for a measurement of , it is anticipated that the precision will improve to a competitive level once a greater number of AGNs are accessible following the upgrade of GRAVITY in the near future.
 From SARM analysis, the black hole masses of the four quasars are also measured with the statistical uncertainty ranging from 0.06 to 0.23 dex, consistent with the correlations between black hole masses and properties of the host bulges.
@@
 2013; Du & Wang 2019).
-Such sizes are too compact to be spatially resolved by modern facilities even for the nearest AGNs.
+Such sizes are too compact to be spatially resolved by modern facilities, even for the nearest AGNs.
 For instance, the Seyfert 1 galaxy NGC 4151 at a distance of 15.8 Mpc (Yuan et al.
@@
 2023), corresponding to an angular size of milli-arcseconds.
-For a long time RM stood out as the most effective method to determine the spatial extent of the BLR (e.g., Kaspi et al.
+For a long time, RM stood out as the most effective method to determine the spatial extent of the BLR (e.g., Kaspi et al.
 2000; Peterson et al.
@@
 2024), provided these have a prominent broad emission line in the band.
-Remarkably, the capabilities of SA and RM techniques are complementary and their combination (hereafter referred to as SARM) conveys a more comprehensive view of the BLR geometry and kinematics.
+Remarkably, the capabilities of SA and RM techniques are complementary and their combination (hereafter referred to as spectroastrometry and reverberation mapping (SARM)) conveys a more comprehensive view of the BLR geometry and kinematics.
 This is reflected in the different dimensions probed by SA and RM.
@@
 The iso-delay surfaces have a paraboloid shape with the focus located at the central ionizing source, typically regarded as a point source coincident with the SMBH.
-The combination of the results from both techniques thus enable a much better overall understanding of the BLR properties as well as precise SMBH mass measurements.
+The combination of the results from both techniques thus enables a much better overall understanding of the BLR properties as well as precise SMBH mass measurements.
 Another important application of the SARM method is that it directly probes geometric distances to AGNs.
@@
 As it is usually difficult to measure both and for the same target, extragalactic objects for which such a geometric method is applicable are exceptionally rare.
-The SARM approach thus offers a new, promising tool for cosmology, especially given ubiquitousness and high brightness of AGNs across the universe.
-Hitherto a sample of seven nearby bright AGNs has been observed by GRAVITY (GRAVITY Collaboration et al.
+The SARM approach thus offers a new, promising tool for cosmology, especially given the ubiquitousness and high brightness of AGNs across the Universe.
+Hitherto, a sample of seven nearby bright AGNs has been observed by GRAVITY (GRAVITY Collaboration et al.
 2018, 2020, 2021a, 2024).
@@
 This radially dependent responsivity of BLR gas has an important implication for SARM analysis.
-As SA probes emissivity weighting whereas RM probes responsivity weighting of the BLR gas, a non-uniform responsivity means that the emissivity- and responsivity-weighted BLR sizes are not the same.
+As SA probes emissivity weighting, whereas RM probes responsivity weighting of the BLR gas, a non-uniform responsivity means that the emissivity- and responsivity-weighted BLR sizes are not the same.
 In this work, we perform SARM analysis on two AGNs investigated by GRAVITY Collaboration et al.
@@
 Section 5 presents discussions on possible systematic uncertainties in the determined distances, followed by the conclusions in Section 6.
-Appendix A defines three types of BLR sizes, Appendix B lists the full set of model parameters and their inferred values, Appendix C shows the best fits to the differential phases of the three baselines with the strongest BLR signals, Appendix D plots the mean and rms spectra of the four AGNs, and Appendix E tests the influence of the upper limit of responsivity on the inferred model parameters.
+Appendix A defines three types of BLR sizes, Appendix B lists the full set of model parameters and their inferred values, Appendix C shows the best fits to the differential phases of the three baselines with the strongest BLR signals, Appendix D plots the mean and rms spectra of the four AGNs, Appendix E tests the influence of the upper limit of responsivity on the inferred model parameters, and Append F compares the obtained time lags from BLR dynamical modeling with those derived from the traditional cross-correlation analysis of the RM data.
 Mrk 509 | PDS 456 | 3C 273 | NGC 3783 | |
 0.0344 | 0.185 | 0.15834 | 0.009730 | |
-GRAVITY Period | Jul. 2021 (2 nights) | Aug. 2018-Jul. 2021 (9 nights) | Jul. 2017-May 2018 (8 nights) | Jan. 2018-Mar. 2020 (6 nights) |
-RM Period | May 2019-Nov. 2022 | May 2020-Oct. 2023 | Nov. 2008-Jul. 2018 | Feb. 2020-Jun. 2020 |
+GRAVITY | Jul. 2021 (2 nights) | Aug. 2018-Jul. 2021 (9 nights) | Jul. 2017-May 2018 (8 nights) | Jan. 2018-Mar. 2020 (6 nights) |
+RM | May 2019-Nov. 2022 | May 2020-Oct. 2023 | Nov. 2008-Jul. 2018 | Feb. 2020-Jun. 2020 |
 2 Observation Data
@@
 (2021a).
-We made additional reduction to the differential phase and line profile data of Mrk 509 and PDS 456 after applying the pipeline following the procedures described in GRAVITY Collaboration et al.
+We made additional reductions to the differential phase and line profile data of Mrk 509 and PDS 456 after applying the pipeline following the procedures described in GRAVITY Collaboration et al.
 (2020).
@@
 The NGC 3783 data were reduced in the same way as described above.
-The 3C 273 data were however reduced differently.
+The 3C 273 data were, however, reduced differently.
 Instead of decomposing the instrumental differential phase signatures into static and variable components, GRAVITY Collaboration et al.
@@
 2017) to yield high-cadence continuum light curves.
-These different sources of data were intercalibrated with the Bayesian software PyCALI111https://github.com/LiyrAstroph/%****␣ms_v5.tex␣Line␣225␣****PyCALI (Li et al.
+These different sources of data were intercalibrated with the Bayesian software PyCALI111The latest code is publicly available at https://github.com/LiyrAstroph/%****␣ms_v5.tex␣Line␣225␣****PyCALI, while the version used in this work is available in Li (2024b).
+(Li et al.
 2014), which uses a multiplicative factor and an additive factor to account for the inhomogeneous apertures and differences in flux calibration.
 For Mrk 509, from the four years of monitoring, the H time delay (around 50 days) can be detected in three years (2019, 2021, and 2022).
-However, when using the full data span, there appears a long-term trend in the H variations, leading to an apparently long time delay component (around 300 days).
+However, when using the full data span, there appears to be a long-term trend in the H variations, leading to an apparently long time delay component (around 300 days).
 The explanation for such a long-term trend is still under debate.
 We therefore only use one-year data for SARM analysis.
-We choose the RM data for year 2022 since it shows the most significant variability.
+We chose the RM data for the year 2022 since it shows the most significant variability.
 We also present analysis with the RM data in 2019 and 2021 and find that the obtained distances are consistent with each other within uncertainties.
@@
 This extends the temporal baseline of the continuum light curve from 137 days to 268 days in the 2020 season and also supplies continuum variation information for other seasons.
-The former is useful to constrain the BLR model and the latter helps to better characterize the continuum variability.
+The former is useful to constrain the BLR model, and the latter helps to better characterize the continuum variability.
 In modeling RM data, we need to subtract the components in the continuum light curves irrelevant to emission-line reverberation.
@@
 2024).
-For NGC 3783, we subtract the host galaxy component () according to the HST image decomposition employed by Bentz et al.
+For NGC 3783, we subtract the host galaxy component () according to the Hubble Space Telescope image decomposition employed by Bentz et al.
 (2021).
@@
 We subtract an averaged jet flux from the continuum light curve according to the decomposition of the jet and accretion disk emissions proposed by Li et al.
-(2020), which derived an averaged jet flux to the total continuum flux fraction of .
+(2020), who derived an averaged jet flux to the total continuum flux fraction of .
 As summarized in Table 1, we note that the GRAVITY and RM observations of all the four objects are undertaken in overlapping periods.
@@
 2022; Li & Wang 2025).
-For the sake of completeness, we briefly describe the major points of SARM analysis below.
+For the sake of completeness, we briefly describe the major points of the SARM analysis below.
 To correct for the cosmological time dilation, we reduce the observation times of the RM light curves by a factor of , where is the redshift of the object.
@@
 3.1 Continuum Light Curve Modeling
-We first need to evaluate the continuum fluxes on a sufficiently well sampled time grid, which is required to integrate reprocessed emission from the whole BLR with a wide range of time delays.
+We first need to evaluate the continuum fluxes on a sufficiently well-sampled time grid, which is required to integrate reprocessed emission from the whole BLR with a wide range of time delays.
 To this end, we employ the damped random walk (DRW) model to delineate the continuum variability, which is a stationary Gaussian process with well-established procedures to interpolate and extrapolate the observed continuum light curve (e.g., Rybicki & Press 1992; Foreman-Mackey et al.
@@
 Specifically, the time length of this period should be equal to or longer than the maximum time delay of the BLR.
-These continuum points plus two hyper-parameters of the DRW process (the damping time scale and variation amplitude) are treated as free parameters and their best-fit values are determined during the subsequent Bayesian inference (see below).
+These continuum points plus two hyperparameters of the DRW process (the damping time scale and variation amplitude) are treated as free parameters, and their best-fit values are determined during the subsequent Bayesian inference (see below).
 3.2 BLR Dynamical Modeling
@@
 Within the opening angle, BLR clouds are distributed uniformly in the azimuthal -direction.
-In the vertical -direction, the angular dispacement of a cloud is assigned as , where the parameter controls the extent to which clouds are clustered over the outer face of the BLR disk and is a random number drawn from a uniform distribution between 0 and 1 (Pancoast et al.
+In the vertical -direction, the angular displacement of a cloud is assigned as , where the parameter controls the extent to which clouds are clustered over the outer face of the BLR disk and is a random number drawn from a uniform distribution between 0 and 1 (Pancoast et al.
 2014).
@@
 From photoionization calculations (Korista & Goad 2004), it is known that the reprocessing coefficient of BLR clouds, denoted by , has a radial dependence.
-The reprocessing coefficient together with the radial distribution of clouds determine the broad-line emissions.
+The reprocessing coefficient , together with the radial distribution of clouds, determines the broad-line emissions.
 To reduce the degrees of freedom, we let be absorbed into the parameterization of the radial distribution and simply set for all clouds.
@@
 The differential phase at velocity is calculated as
-where is the flux ratio of the emission line relative to the underlying continuum, is coordinate of the baseline, is the emissivity-weighted offset of the BLR projected onto the sky, is the offset of the continuum emission, and is the angular-diameter distance.
+where is the flux ratio of the emission line relative to the underlying continuum, is the coordinate of the baseline, is the emissivity-weighted offset of the BLR projected onto the sky, is the offset of the continuum emission, and is the angular-diameter distance.
 The offset is determined by summing over all clouds as
@@
 The SA data include the profile and differential phases of the broad Pa or Br line.
-The corresponding likelihood function is written
+The corresponding likelihood function is written as
 where the subscript represents -th velocity bin, the superscript represents the modeled quantity, and are the differential phase and uncertainty, and and are the line flux density and uncertainty, respectively.
@@
 2011) to explore the parameter space and generate Markov chains, which has been shown to be effective at exploring multimodal distributions and strong correlations between parameters.
-We have incorporated the above BLR dynamical modeling into our previously developed package BRAINS333The latest code is publicly available at https://github.com/LiyrAstroph/BRAINS, while the version used in this work is available at https://doi.org/10.5281/zenodo.10700132.
+We have incorporated the above BLR dynamical modeling into our previously developed package BRAINS333The latest code is publicly available at https://github.com/LiyrAstroph/BRAINS, while the version used in this work is available in Li (2024a).
 (Li et al.
-2013, 2018), which utilizes the DNS code CDNest444The latest code is publicly available at https://github.com/LiyrAstroph/CDNest, while the version used in this work is available at https://doi.org/10.5281/zenodo.10700132..
+2013, 2018), which utilizes the DNS code CDNest444The latest code is publicly available in https://github.com/LiyrAstroph/CDNest, while the version used in this work is available in Li (2023)..
 Both BRAINS and CDNest are written in C Language with the standardized message passing interface and are thereby portable to a wide range of supercomputer clusters.
@@
 The analysis reproduces only the averaged variation trend of the H light curve.
-As a result, there appear quite a few data points deviating from the best reconstruction.
+As a result, there appear a number of data points deviating from the best reconstruction.
 In the averaged differential phase curves (with the continuum phase subtracted) plotted in the second panel of the top left row of Figures 1-4, 3C 273 and NGC 3783 display an approximately S-shaped signal typical for Keplerian rotation, which is characterized by a crest and trough respectively on either side of the center wavelength of the emission line.
@@
 In the present dynamical modeling, the latter can be explained by the presence of a fraction of BLR clouds with inflowing/outflowing motion plus anisotropic emissions.
+For PDS 456, there appears an extra red-wing component around 2.24 m, which is not reproduced by our current model.
+We emphasize that the plotted differential phase curve indeed represents the averaged data from the three baselines with the strongest BLR phase signals (see Appendix C).
+Notably, this component does not appear across all baselines, possibly implying that it might be caused by some unknown measurement noise or a real BLR component at a special location and velocity that only produces signals along certain baselines.
+With the available quantity and quality of data we cannot discriminate between these two possibilities.
+Future higher-quality data will facilitate either a more advanced BLR modeling or a confirmation that we have here a transient observing feature if the extra red-wing component is real.
 4.2 The Inferred Radial-dependent Responsivity
@@
 Our improved SARM analysis naturally accounts for the differences in these two BLR sizes and largely alleviates the increased systematic biases in distance measurements.
-As mentioned in the preceding section, we find that 3C 273 shows strong radial dependence of the responsivity, while this dependence is weak in NGC 3783 (see Figure 5).
+As mentioned in the preceding section, we find that 3C 273 shows a strong radial dependence of the responsivity, while this dependence is weak in NGC 3783 (see Figure 5).
 These results remarkably coincide with the line widths in the mean and rms spectra of the two objects (see Appendix D).
@@
 If simply scaling results with the above estimated differences in the BLR sizes, the geometric distance will increase by roughly 0.07, 0.19, 0.02, and 0.14 dex for PDS 456, 3C 273, Mrk 509, and NGC 3783, respectively.
-However, we note that a precise evaluation of the influences to the distance measurements requires comprehensive modeling of the H and Pa/Br lines simultaneously, which is beyond the scope of the present work.
+However, we note that a precise evaluation of the influences on the distance measurements requires comprehensive modeling of the H and Pa/Br lines simultaneously, which is beyond the scope of the present work.
 Nevertheless, our estimate of the Hubble constant remains within a reasonable range, indicating that the discrepancy arising from different emission lines measured by SA and RM is moderate.
-In future, this issue can be surmounted by near-infrared RM campaigns observing the same broad emission line as SA or a joint analysis of SA and velocity-resolved RM.
+In the future, this issue can be surmounted by near-infrared RM campaigns observing the same broad emission line as SA or a joint analysis of SA and velocity-resolved RM.
 In the latter approach, one no longer needs to assume the same region for the H and Pa/Br lines but only let the two lines share the same central SMBH (Li et al.
@@
 2024).
-We choose the 2022 RM data for SARM analysis given that this year marked the largest variability.
-In Figure 11, we illustrate a comparison of the obtained BLR size, SMBH mass, and angular-diameter distance between years 2019, 2021, and 2022.
+We chose the 2022 RM data for SARM analysis given that this year marked the largest variability.
+In Figure 11, we illustrate a comparison of the obtained BLR size, SMBH mass, and angular-diameter distance between the years 2019, 2021, and 2022.
 As observed, the best estimates are consistent with each other within uncertainties, although the BLR size in 2021 shows a slightly larger median value.
@@
 These values are in good agreement with known relations between black hole masses and stellar velocity dispersions or dynamical masses of the host bulges (though for PDS 456 without relevant measurements of the host galaxy).
-Although our estimated Hubble constant still carries a large uncertainty, the forthcoming upgraded instrument GRAVITY+ will be able to observe more AGNs (up to a few hundred) with higher-quality differential phase measurements (GRAVITY+ Collaboration et al.
+The currently observed AGNs by GRAVITY remain rare due to the strict magnitude limit (; GRAVITY Collaboration et al.
+2017).
+However, the forthcoming upgraded instrument GRAVITY+ will enable phase reference on a target with using the same observation mode (“on-axis”555Alternatively, with the “off-axis” mode, fainter AGNs () are observable provided there is a bright star () within a 30″ separation serving as the phase reference.
+However, due to reduced photon fluxes, these AGNs require extended exposure times to achieve an adequate signal-to-noise ratio for BLR differential phases observations.) as the targets reported here.
+This will permit interferometric observations of a much larger number of AGNs (up to a few hundred; GRAVITY+ Collaboration et al.
+2022) and thereby can significantly expand the AGN sample for SARM analysis.
+The improved sensitivity of GRAVITY+ can also facilitate higher-quality differential phase measurements.
+Given the feasibility of RM monitoring, those AGNs observable with the on-axis mode are vital for geometric distance determinations.
+In particular, the Pa line, redshifted into the band for , is generally about 10 times brighter than the Br line for .
+Therefore, the Pa AGN sample is preferred for future SARM observations with GRAVITY+.
+In addition, a new visitor interferometry instrument BIFROST at the and bands on VLTI will come into operation within the next few years (Kraus et al.
 2022).
-A sizeable RM campaign coordinated with GRAVITY+ interferometric observations will enable the SARM approach to achieve a competitive precision for the Hubble constant.
+It is capable of resolving the prominent Pa and Pa lines in low- AGNs ().
+Moreover, advancements in fringe-tracking (FT) methodologies (such as the hierarchical FT architecture; see Petrov et al.
+2024) might enhance the FT limiting magnitude and thus further expand the accessible AGN sample size.
+Taken together, we would be able to observe sufficient AGN targets for a statistically relevant combination of high-quality SA with RM.
+The high-quality SA would also allow to eliminate or better model those targets potentially exhibiting unusual features like PDS 456.
+Although our currently estimated Hubble constant still carries a large uncertainty, a sizable RM campaign coordinated with near-infrared interferometry will enable the SARM approach to achieve a competitive precision for the Hubble constant.
 In light of the geometric nature of the measured distance, the SARM approach will ultimately offer a promising cosmic probe, independent of all other cosmic tools.
 Acknowledgements
-We thank Misty Bentz for kindly sharing the RM light-curve data of NGC 3783 and Hannah Übler for useful comments on the manuscript.
-YRL acknowledges financial support from the National Key R&D Program of China (2021YFA1600404, 2023YFA1607904), the National Natural Science Foundation of China (12273041), and from the Youth Innovation Promotion Association CAS.
+We thank Misty Bentz for kindly sharing the RM light-curve data of NGC 3783 and Hannah Übler and the anonymous referee for useful comments on the manuscript.
+YRL acknowledges financial support from the National Key R&D Program of China (2021YFA1600404, 2023YFA1607904), the National Natural Science Foundation of China (12273041), the China-Chile Joint Research Fund (CCJRF2310), and from the Youth Innovation Promotion Association CAS.
+JS is supported by the Fundamental Research Funds for the Central Universities, Peking University (7100604896) and the China Manned Space Program (CMS-CSST-2025-A09).
 JMW acknowledges financial support from the National Natural Science Foundation of China (11991050, 12333003).
-LCH was supported by the National Key R&D Program of China (2022YFF0503401), the National Science Foundation of China (11991052, 12233001), and the China Manned Space Project (CMS-CSST-2021-A04, CMS-CSST-2021-A06).
-We thank the support of the staff of the Lijiang 2.4 m telescope and of the CAHA 2.2m telescope.
-This work is based on observations collected at the Centro Astronómico Hispanoen Andalucía (CAHA) at Calar Alto, operated jointly by the Andalusian Universities and the Instituto de Astrofísica de Andalucía (CSIC).
+LCH is supported by the National Key R&D Program of China (2022YFF0503401), the National Science Foundation of China (11991052, 12233001), and the China Manned Space Project (CMS-CSST-2021-A04, CMS-CSST-2021-A06).
+P.D.
+is supported by the National Key R&D Program of China (2023YFA1607903) and the China Manned Space Project (CMS-CSST-2025-A07).
+R.G.P.
+acknowledges the support of the French Agence Nationale de la Recherche (ANR) through the grant “AGN MELBa” ANR-21-CE31-0011.
+We thank the support of the staff of the Lijiang 2.4 m telescope and of the the Centro Astronómico Hispanoen Andalucía (CAHA) 2.2m telescope.
+This work is based on observations collected at CAHA at Calar Alto, operated jointly by the Andalusian Universities and the Instituto de Astrofísica de Andalucía (CSIC).
 This work has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.
@@
 Nevertheless, the differences from those of and are about 0.1 dex, still within 1 uncertainties, implying that the influence of on the distance measurement is not significant.
-A BLR dynamical model with an inclusion of photoionization models will relax the need of present simple parameterization for the responsivity and help further improve SARM analysis.
+A BLR dynamical model with an inclusion of photoionization models will relax the need for the present simple parameterization for the responsivity and help further improve SARM analysis.
 This is beyond the scope of this paper and will be deferred to a future work.
+Appendix F Comparison with Time Lags from the Cross-correlation Functions
+Given a BLR model, we can directly calculate the centroid time lag of the responsivity-weighted transfer function using Equation (A1).
+In practice, the transfer function usually has a long tail, arising from the responses of BLR clouds at large distances.
+This long tail causes the centroid time lag of the transfer function to be longer than that from the traditional cross-correlation function (CCF) of the light curves.
+The latter is conventionally derived by the centroid of the CCF above 80% of the peak value (e.g., Gaskell & Peterson 1987; Peterson et al.
+2004).
+Therefore, to ensure a meaningful comparison, we calculate the time lag from BLR modeling as follows.
+We firstly generate mock datasets based on the reconstructed light curves from the BLR model by adopting the same measurement errors and sampling cadences as the observed data.
+We then use the same CCF method to calculate the time lags of these mock datasets.
+Figure 15 shows the comparison.
+The time lags from the two approaches are consistent, except for PDS 456, which shows a relatively shorter time lag from BLR modeling (but still consistent at confidence level).
+We ascribe this discrepancy to the influence of fine variation patterns in the H light curve of PDS 456 (see Figure 2), which appear to lack counterparts in the continuum variations and thereby cannot be reproduced by RM modeling.
 Parameter | Prior | Range | Note |
 (ld) | LogUniform | (1, /2) | Emissivity-weighted mean BLR size |
-Uniform | (0, 1) | Fraction of the inner BLR edge in unit of | |
+Uniform | (0, 1) | Fraction of the inner BLR edge in units of | |
 Uniform | (0, 2) | Shape parameter of radial distribution of BLR clouds | |
@@
 — denotes the time span of the emission line light curve.
-“LogUniform” denotes a uniform prior distribution for the logarithm of the parameter and “CosUniform” denote a uniform prior distribution for the cosine of the parameter.
-Parameter | Mrk 509 | PDS 456 | 3C 273 | NGC 3783 |
-PA (∘ E of N) | ||||
+“LogUniform” denotes a uniform prior distribution for the logarithm of the parameter, and “CosUniform” denotes a uniform prior distribution for the cosine of the parameter.
+Parameter | Mrk 509 | PDS 456aaThe continuum phase offset of PDS 456 is set to zero. | 3C 273 | NGC 3783 |
+PA (∘ E of N)bbNote that the position angle (PA) has a 180∘ ambiguity. | ||||
 (ld) | 0 | |||